OpenStudy (anonymous):

Calculate the moment of inertia of a bicycle wheel 66.7 cm in diameter. The rim and tire have a combined mass of 1.25 kg. The mass of the hub can be ignored (why?). I'm not sure how to start or set this up. I'm sure that I have to use the formula I=Emr^2 Would I put 1.25kg as "m" and convert 66.7cm to meters and divide that answer by 2 to make it a radius? I'm completely guessing here!

OpenStudy (mos1635):

I=M*r^2 E=???

OpenStudy (anonymous):

use I=mr^2